6.5) Other things equal, and assuming efficient markets, if a Honda Accord costs $21,375 in the U.S. then at an exchange rateplease answer all parts of question 6.
Answer #1

6.5) Honda cost in US = $21,375

Exchange Rate: $/€ = 1.23

Therefore Home cost in France = 21,375/1.23 = €17,378.05

6.6) let the inflation in Canada = 4%

Then The inflation in USA is 3% less than Canada = 1%

One year ago $/C$ = 1. (S​​​​​​0​​​​​)

The current exchange rate (S​​​​​​1) = ?

As per PPP

S1/S0 = (1+inflation of USD)/(1+inflation of Canada)

S​​​​1​​​​​÷1 = (1+0.01)/(1+0.04)

S1 = 1.01/1.04

S1 = 0.97

The current exchange rate (S​​​​​​1): $/C$= 0.97

6.7)

Last year price is US = 150USD

This year Price in US = 155USD

% change in price = 155-150/150 = 3.33333%

Last year Exchange rate : USD/Euro = 1.25

This year exchange rate : USD/Euro = 1.30

% change in Exchange rates = 1.3-1.25/1.25 = 4%

Exchange rate pass through =

% change in Price /% change in Exchange rate.

3.3333/4 = 0.83325.

This means: For every 1% increase in the exchange rate, there has been a .83325% increase in the price of the DVD player.

6.8) Forward exchange rate after 90 days (£/$) = 0.701035

6.8 t/$0.6993()usp 3/ As per Intest Pate partly theory 4/- 1+tSA B 1+t88 a we have asuned yeae con tains 260dayg 1+(10C0.04 F